# GD-VAEs: Geometric Dynamic Variational Autoencoders for Learning Nonlinear Dynamics and Dimension Reductions

### v3

@@ Introduction
The general problem of learning dynamical models from a time series of observations has a long history spanning many fields [59; 76; 18; 38]. This includes the fields of dynamical systems [10; 76; 78; 55; 58; 60; 34; 21; 25], control [11; 59; 70; 74], statistics [1; 56; 28], and machine learning [18; 38; 51; 68; 5; 86; 52]. Many of the most successful and widely-used approaches rely on assumptions on the model structure, most commonly, that a time-invariant linear dynamical systems (LDS) provides a good local approximation or that the noise is Gaussian. These include the Kalman Filter and extensions [43; 24; 30; 82; 84], Proper Orthogonal Decomposition (POD) [15; 57], and more recently Dynamic Mode Decomposition (DMD) [74; 49; 81] and Koopman Operator approaches [58; 22; 47]. Methods for learning nonlinear dynamics include the NARX and NOE approaches with function approximators based on neural networks and other models classes [59; 76], sparse symbolic dictionary methods that are linear-in-parameters using LASSO [71; 79] such as SINDy [11; 75; 76], and dynamic Bayesian networks (DBNs), such as Hidden Markov Chains (HMMs) and Hidden-Physics Models  [68; 62; 73; 7; 48; 28].
@@
Many variants of autoencoders have been developed for making predictions of sequential data. This includes those based on Recurrent Neural Networks (RNNs) with LSTMs and GRUs [37; 31; 19]. Approaches for incorporating topological information into latent variable representations include the early works by Kohonen on Self-Organizing Maps (SOMs) [46] and Bishop on Generative Topographical Maps (GTMs) based on density networks providing a generative approach [8]. While general RNNs provide a rich approximation class for sequential data, they pose for dynamical systems challenges for interpretability and for training to obtain predictions stable over many steps. Autoencoders have also been combined with symbolic dictionary learning for latent dynamics in [14] providing some advantages for interpretability and robustness. Dictionary methods however require specification in advance of sufficiently expressive libraries of functions. Neural networks incorporating physical information have also been developed, where some of the methods use regularizations introduced during training to enhance stability [61; 3; 80; 51; 80; 4; 26]. The work of [20] considers methods for processing of speech and handwriting by investigating RNNs combined with VAEs to obtain more robust sequential models.
@@
In our work, we learn dynamical models building on the VAE framework to train probabilistic encoders and decoders between general manifold latent spaces. This provides additional regularizations and constraints to help promote parsimoniousness, disentanglement of features, robustness, and interpretability. Prior VAE methods used for dynamical systems include [33; 63; 52; 29; 16; 63; 69]. These works use primarily Euclidean latent spaces and consider applications such as human motion capture and ODE systems. More recently, VAE methods using non-Euclidean latent spaces include [41; 42; 17; 52; 2]. These incorporate the role of geometry by augmenting the prior distribution on a Euclidean latent space to bias encodings toward a manifold. In the recent works [65; 23; 27], explicit projection procedures are introduced to map analytically or to sample through random walks an embedded manifold. These works primarily consider geometries based on $SO(3)$, spheres, tori, and cylinders. We develop in our work more explicit procedures that can be used with back-propagation for more general geometries, including non-orientable manifolds. We also discuss related methods in our proceedings paper [53].
@@ 1 Learning Nonlinear Dynamics with Variational Autoencoders (VAEs)
Figure 1: Learning Nonlinear Dynamics with Encoders and Decoders. Data-driven methods are developed for learning robust models to predict from $u(x,t)$ the non-linear evolution to future states $u(x,t+\tau)$ for PDEs and other dynamical systems. Probabilistic autoencoders are utilized to learn representations $\mathbf{z}$ of $u(x,t)$ in reduced dimensional latent spaces with prescribed geometric and topological properties. Our trainable models make predictions using learnable maps that (i) encode an input $u(x,t)\in\mathcal{U}$ as $\mathbf{z}(t)$ in latent space (left), (ii) evolve the representation $\mathbf{z}(t)\rightarrow\mathbf{z}(t+\tau)$ (middle), (iii) decode the representation $\mathbf{z}(t+\tau)$ to predict $\hat{u}(x,t+\tau)$ (right).
We learn VAE predictors using a Maximum Likelihood Estimation (MLE) approach. For dynamic predictions $\mathbf{X}\rightarrow\mathbf{x}$, we use a Log Likelihood (LL) which can be expressed as
@@
The $\mathcal{J}_{LC}$ gives the Log-Conditional (LC) Likelihood and $\mathcal{J}_{LX}$ gives the marginal Log Likelihood of $\mathbf{X}$. For predicting the dynamics of $u(t)$, we discretize in time. We sample trajectories as $\{u(s_{k})\}_{k=0}^{N}$, where $s_{k}=t_{0}+k\tau$ and $\tau$ is the sampling time-scale. We learn representations for the discretized dynamics $\mathbf{X}\rightarrow\mathbf{x}$ by letting $\mathbf{X}=u(s_{k})$ and $\mathbf{x}=u(s_{k+1})$. We base $p_{\theta}$ on the autoencoder framework in Figure 1 and 2. More specifically, we use variational inference to approximate the LL by the Evidence Lower Bound (ELBO) [9; 45]. In particular, we use the concavity of $\log(\cdot)$ and Jensen’s Inequality for $\log\mathbb{E}[(\cdot)]\geq\mathbb{E}[\log(\cdot)]$. This yields the lower-bound
@@
To simplify the notation for the expressions above, we have suppressed dependence on the parameters $\theta$. We assume here the considered probability distributions factor to have an autoencoding-like property so that $p(\mathbf{X},\mathbf{x},\mathbf{z}):=p(\mathbf{x}|\mathbf{z})p(\mathbf{z}|\mathbf{X})p(\mathbf{X})$. This ensures the encoding $\mathbf{z}$ captures all the relevant information to generate $\mathbf{x}$, so that $p(\mathbf{x}|\mathbf{z},\mathbf{X})=p(\mathbf{x}|\mathbf{z})$. The $q(\mathbf{x}|\mathbf{z})$ gives the distribution for a probabilistic encoder map from $\mathbf{x}$ to $\mathbf{z}$. The $p(\mathbf{x}|\mathbf{z})$ gives a probabilistic decoder map that reconstructs $\mathbf{x}$ from $\mathbf{z}$.
The encoding and decoding distributions in the variational approximations will often be taken to be learnable Gaussians of the form $p_{\theta}(\mathbf{x}|\mathbf{z})=\mathcal{N}(\mathbf{x};\boldsymbol{\mu}_{d}(\mathbf{z};\theta),\boldsymbol{\Sigma}_{d}(\theta))$ and $q_{\theta}(\mathbf{z}|\mathbf{X})=\mathcal{N}(\mathbf{z};\boldsymbol{\mu}_{e}(\mathbf{X};\theta),\boldsymbol{\Sigma}_{e}(\theta))$. The $\mathcal{N}$ denotes the probability density of the multi-variate Gaussian with mean $\boldsymbol{\mu}(\mathbf{z};\theta)$ and covariance $\boldsymbol{\Sigma}(\theta)$. In the case of no noise $\boldsymbol{\Sigma}_{e}=0$, the $\boldsymbol{\mu}_{e}(\mathbf{X})$ would serve as a deterministic encoding map of $\mathbf{x}$ to obtain $\mathbf{z}$. When there is noise $\boldsymbol{\Sigma}_{e}\neq 0$, $\boldsymbol{\mu}_{e}$ becomes the center location of the Gaussian in latent space and is the most probable encoding of $\mathbf{x}$ to obtain $\mathbf{z}$, see Figure 2 (lower-left). This holds similarly for the probabilistic decoder map given by $\boldsymbol{\mu}_{d}$ and $\boldsymbol{\Sigma}_{d}$. For the decoder, the log probability gives a weighted $\ell^{2}$-norm penalizing errors in reconstructing $\mathbf{x}$, since for a Gaussian $-\log(p(\mathbf{x}|\mathbf{z}))=\frac{1}{2}\|\mathbf{x}-\boldsymbol{\mu}(\mathbf{z})\|_{\Sigma}^{2}+c_{0}$. Here, the $c_{0}$ is a constant independent of $\mathbf{x}$ and
@@
It is convenient to reformulate the maximization of the terms of the lower-bound in equation 1 to a minimization problem with loss function $\mathcal{L}_{LL}=-\mathcal{J}_{LL}$. The optimization then can be interpreted as minimizing the reconstruction losses
$\mathcal{L}_{RE}$ $=$ $-\mathcal{J}_{RE}=-\mathbb{E}_{q(\mathbf{z}|\mathbf{X})}\left[\log\left(p(\mathbf{x}|\mathbf{z})\right)\right]$ (4)
@@
subject to further regularization based on reducing the Kullback-Lieber (KL)-Divergence between $q(\mathbf{z}|\mathbf{X})$ and $p(\mathbf{z})$,
@@
$\mathcal{L}_{RE}^{(i)}$ $=$ $-E_{\mathfrak{q}_{\theta_{e}}(\mathbf{z}|\mathbf{X}^{(i)})}\left[\log\mathfrak{p}_{\theta_{d}}(\mathbf{x}^{(i)}|\mathbf{z}^{\prime})\right],\;\;\mathcal{L}_{KL}^{(i)}=\beta\mathcal{D}_{KL}\left(\mathfrak{q}_{\theta_{e}}(\mathbf{z}|\mathbf{X}^{(i)})\,\|\,\tilde{\mathfrak{p}}_{\theta_{d}}(\mathbf{z})\right),$
@@
The $\mathfrak{q}_{\theta_{e}}$ denotes the distribution for the probabilistic encoding map, $\mathfrak{p}_{\theta_{d}}$ denotes the distribution of the probabilistic decoding map, and $\tilde{\mathfrak{p}}_{\theta_{d}}$ the prior distribution used for regularizing the organization of the latent-space representation $\bar{\mathbf{z}}$. This is trained using batches of $m$ observation samples $\{(\mathbf{X}^{(i)},\mathbf{x}^{(i)})\}_{i=1}^{m}$. For a dynamical system with a collection of sampled trajectories $\{u^{k}(s)\}_{k}$, we have that $\mathbf{X}^{(i)}$ samples the initial state $u^{k_{i}}(t_{i})$ and $\mathbf{x}^{(i)}$ samples the future state $u^{k_{i}}(t_{i}+\tau)$, where $k_{i}$ gives the trajectory index and $t_{i}$ the sampled initial time.
The loss function in equation 7 provides a regularized form of MLE. The three terms comprising $\mathcal{L}^{B}$ can be interpreted as follows (i) $\mathcal{L}_{RR}$ for the log likelihood of reconstructing samples, (ii) $\mathcal{L}_{RE}$ is the log likelihood of correctly predicting samples after a single time step, and (iii) $\mathcal{L}_{KL}$ is a regularization term based on the Kullback-Liebler (KL) Divergence between the encoding distribution and latent space prior distribution.
@@
$\begin{array}[]{llllll}\mathbf{z}&\sim&\mathfrak{q}_{\theta_{e}}(\mathbf{z}|\mathbf{X}),&\mathbf{z}&=&\boldsymbol{\mu}_{e}(\mathbf{X};\theta_{e})+\boldsymbol{\eta}_{e}\\ \mathbf{z}^{\prime}&\sim&\mathfrak{p}_{\theta_{\ell}}(\mathbf{z}^{\prime}|\mathbf{z}),&\mathbf{z}^{\prime}&=&f_{\theta_{\ell}}(\mathbf{z})\\ \mathbf{x}&\sim&\mathfrak{p}_{\theta_{d}}(\mathbf{x}|\mathbf{z}^{\prime}),&\mathbf{x}&=&\boldsymbol{\mu}^{\prime}_{d}(\mathbf{z}^{\prime};\theta_{d})+\boldsymbol{\eta}_{d}^{\prime}\\ \mathbf{X}&\sim&\mathfrak{p}_{\theta_{d}}(\mathbf{X}|\mathbf{z}),&\mathbf{X}&=&\boldsymbol{\mu}_{d}(\mathbf{z};\theta_{d})++\boldsymbol{\eta}_{d}.\end{array}$ (15)

### v4

@@ Introduction
The general problem of learning dynamical models from a time series of observations has a long history spanning many fields [59; 78; 18; 38]. This includes the fields of dynamical systems [10; 78; 81; 55; 58; 60; 34; 21; 25], control [11; 59; 72; 76], statistics [1; 56; 28], and machine learning [18; 38; 51; 70; 5; 89; 52]. Many of the most successful and widely-used approaches rely on assumptions on the model structure, most commonly, that a time-invariant linear dynamical systems (LDS) provides a good local approximation or that the noise is Gaussian. These include the Kalman Filter and extensions [43; 24; 30; 85; 87], Proper Orthogonal Decomposition (POD) [15; 57], and more recently Dynamic Mode Decomposition (DMD) [76; 49; 84] and Koopman Operator approaches [58; 22; 47]. Methods for learning nonlinear dynamics include the NARX and NOE approaches with function approximators based on neural networks and other models classes [59; 78], sparse symbolic dictionary methods that are linear-in-parameters using LASSO [73; 82] such as SINDy [11; 77; 78], and dynamic Bayesian networks (DBNs), such as Hidden Markov Chains (HMMs) and Hidden-Physics Models  [70; 62; 75; 7; 48; 28].
@@
Many variants of autoencoders have been developed for making predictions of sequential data. This includes those based on Recurrent Neural Networks (RNNs) with LSTMs and GRUs [37; 31; 19]. Approaches for incorporating topological information into latent variable representations include the early works by Kohonen on Self-Organizing Maps (SOMs) [46] and Bishop on Generative Topographical Maps (GTMs) based on density networks providing a generative approach [8]. While general RNNs provide a rich approximation class for sequential data, they pose for dynamical systems challenges for interpretability and for training to obtain predictions stable over many steps. Autoencoders have also been combined with symbolic dictionary learning for latent dynamics in [14] providing some advantages for interpretability and robustness. Dictionary methods however require specification in advance of sufficiently expressive libraries of functions. Neural networks incorporating physical information have also been developed, where some of the methods use regularizations introduced during training to enhance stability [61; 3; 83; 79; 51; 83; 4; 26]. The work of [20] considers methods for processing of speech and handwriting by investigating RNNs combined with VAEs to obtain more robust sequential models.
@@
In our work, we learn dynamical models building on the VAE framework to train probabilistic encoders and decoders between general manifold latent spaces. This provides additional regularizations and constraints to help promote parsimoniousness, disentanglement of features, robustness, and interpretability. Prior VAE methods used for dynamical systems include [33; 63; 52; 29; 16; 63; 71]. These works use primarily Euclidean latent spaces and consider applications such as human motion capture and ODE systems. More recently, VAE methods using non-Euclidean latent spaces include [41; 42; 17; 52; 2]. These incorporate the role of geometry by augmenting the prior distribution on a Euclidean latent space to bias encodings toward a manifold. In the recent works [65; 23; 27], explicit projection procedures are introduced to map analytically or to sample through random walks an embedded manifold. These works primarily consider geometries based on $SO(3)$, spheres, tori, and cylinders. We develop in our work more explicit procedures that can be used with back-propagation for more general geometries, including non-orientable manifolds. We discuss related methods in our proceedings paper [53]. We also discuss related methods for handling geometry using neural operator approaches in our recent papers [69; 68] and stochastic systems in [79].
@@ 1 Learning Nonlinear Dynamics with Variational Autoencoders (VAEs)
Figure 1: Learning Nonlinear Dynamics with Encoders and Decoders. Data-driven methods are developed for learning models to predict from $u(x,t)$ the non-linear evolution to future states $u(x,t+\tau)$ for PDEs and other dynamical systems. Probabilistic autoencoders are utilized that learn representations $\mathbf{z}$ of $u(x,t)$ in reduced dimensional latent spaces with prescribed geometric and topological properties. The trainable models make predictions using learnable maps that (i) encode an input $u(x,t)\in\mathcal{U}$ as $\mathbf{z}(t)$ in latent space (left), (ii) evolve the representation $\mathbf{z}(t)\rightarrow\mathbf{z}(t+\tau)$ (middle), (iii) decode the representation $\mathbf{z}(t+\tau)$ to predict $\hat{u}(x,t+\tau)$ (right).
We learn VAE predictors using a Maximum Likelihood Estimation (MLE) approach. For dynamic predictions $\mathbf{X}\rightarrow\mathbf{x}$, we use a Log Likelihood (LL) which can be expressed using the factorization $p_{\theta}(\mathbf{X},\mathbf{x})=p_{\theta}(\mathbf{x}|\mathbf{X}))p_{\theta}(\mathbf{X})$ as
@@
The $\mathcal{J}_{LC}$ gives the Log-Conditional (LC) Likelihood and $\mathcal{J}_{LX}$ gives the marginal Log Likelihood of $\mathbf{X}$. For predicting the dynamics of $u(t)$, we discretize in time. We sample trajectories as $\{u(s_{k})\}_{k=0}^{N}$, where $s_{k}=t_{0}+k\tau$ and $\tau$ is the sampling time-scale. We learn representations for the discretized dynamics $\mathbf{X}\rightarrow\mathbf{x}$ by letting $\mathbf{X}=u(s_{k})$ and $\mathbf{x}=u(s_{k+1})$ with $p_{\theta}$ based on the dynamic autoencoder framework in Figure 1 and 2. We use variational inference to approximate the LL by the Evidence Lower Bound (ELBO) [9; 45]. We use the concavity of $\log(\cdot)$ and Jensen’s Inequality for $\log\mathbb{E}[(\cdot)]\geq\mathbb{E}[\log(\cdot)]$. This yields the lower-bound
@@
Maximizing this lower bound corresponds to minimizing the loss $\mathcal{L}^{B}=-\mathcal{J}_{RE}-\mathcal{J}_{RR}-\mathcal{J}_{KL}$. We treat the $\mathcal{J}_{RR}$ and $\mathcal{J}_{KL}$ as regularization terms for the prediction and introduce additional adjustable parameters $\gamma,\beta$. We use the regularized loss function
$\mathcal{L}^{B}=-\mathcal{J}_{RE}-\gamma\mathcal{J}_{RR}-\beta\mathcal{J}_{KL}.$ (3)
To simplify the notation for the expressions above, we have suppressed dependence on the distribution parameters $\theta$. We assume here the considered probability distributions factor to have an autoencoding-like property so that $p(\mathbf{X},\mathbf{x},\mathbf{z}):=p(\mathbf{x}|\mathbf{z})p(\mathbf{z}|\mathbf{X})p(\mathbf{X})$. This ensures the encoding $\mathbf{z}$ captures all the relevant information to generate $\mathbf{x}$, so that $p(\mathbf{x}|\mathbf{z},\mathbf{X})=p(\mathbf{x}|\mathbf{z})$. The $q(\mathbf{x}|\mathbf{z})$ gives the distribution for a probabilistic encoder map from $\mathbf{x}$ to $\mathbf{z}$. The $p(\mathbf{x}|\mathbf{z})$ gives a probabilistic decoder map that reconstructs $\mathbf{x}$ from $\mathbf{z}$.
We remark that the dynamic VAE can also be viewed abstractly as an autoencoder for reconstructing $(\mathbf{X},\mathbf{x})$ with constraints on the distributions for encoding and decoding. In particular, the factorization $p(\mathbf{X},\mathbf{x},\mathbf{z})=p(\mathbf{x}|\mathbf{z})p(\mathbf{z}|\mathbf{X})p(\mathbf{X})$ corresponds to $\mathbf{z}$ making $\mathbf{X}$ and $\mathbf{x}$ independent as a data processing Markov chain $\mathbf{X}\rightarrow\mathbf{z}\rightarrow\mathbf{x}$. The encoder distribution is constrained only to have a dependence on $\mathbf{X}$ as $q(\mathbf{z}|\mathbf{X},\mathbf{x})=q(\mathbf{z}|\mathbf{X})$. In the limit $\gamma\rightarrow 0$, which would remove the initial condition reconstruction regularization $\mathcal{J}_{RR}$, the $\mathbf{X}$ would be encoded only to reconstruct $\mathbf{x}$. We discuss the consequences of this and other parameter choices in our studies below.
The encoding and decoding distributions in the variational approximations will be taken to be learnable Gaussians of the form $p_{\theta}(\mathbf{x}|\mathbf{z})=\mathcal{N}(\mathbf{x};\boldsymbol{\mu}_{d}(\mathbf{z};\theta),\boldsymbol{\Sigma}_{d}(\theta))$ and $q_{\theta}(\mathbf{z}|\mathbf{X})=\mathcal{N}(\mathbf{z};\boldsymbol{\mu}_{e}(\mathbf{X};\theta),\boldsymbol{\Sigma}_{e}(\theta))$. The $\mathcal{N}$ denotes the probability density of the multi-variate Gaussian with mean $\boldsymbol{\mu}(\mathbf{z};\theta)$ and covariance $\boldsymbol{\Sigma}(\theta)$. In the case of no noise $\boldsymbol{\Sigma}_{e}=0$, the $\boldsymbol{\mu}_{e}(\mathbf{X})$ would serve as a deterministic encoding map of $\mathbf{x}$ to obtain $\mathbf{z}$. When there is noise $\boldsymbol{\Sigma}_{e}\neq 0$, $\boldsymbol{\mu}_{e}$ becomes the center location of the Gaussian in latent space and is the most probable encoding of $\mathbf{x}$ to obtain $\mathbf{z}$, see Figure 2 (lower-left). This holds similarly for the probabilistic decoder map given by $\boldsymbol{\mu}_{d}$ and $\boldsymbol{\Sigma}_{d}$. For the decoder, the log probability gives a weighted $\ell^{2}$-norm penalizing errors in reconstructing $\mathbf{x}$, since for a Gaussian $-\log(p(\mathbf{x}|\mathbf{z}))=\frac{1}{2}\|\mathbf{x}-\boldsymbol{\mu}(\mathbf{z})\|_{\Sigma}^{2}+c_{0}$. Here, the $c_{0}$ is a constant independent of $\mathbf{x}$ and
@@
As mentioned above, we reformulate the maximization of the lower-bound in equation 1 as a minimization of the regularized loss $\mathcal{L}^{B}$. The optimization then can be interpreted as minimizing the reconstruction loss for the predicted future state $\mathbf{x}$ subject to additional regularizations
$\mathcal{L}_{RE}$ $=$ $-\mathcal{J}_{RE}=-\mathbb{E}_{q(\mathbf{z}|\mathbf{X})}\left[\log\left(p(\mathbf{x}|\mathbf{z})\right)\right].$ (5)
This includes regularization by reconstruction of the initial conditions $\mathbf{X}$
@@
and regularization based on reducing the Kullback-Leiber (KL)-Divergence between $q(\mathbf{z}|\mathbf{X})$ and the prior distribution $p(\mathbf{z})$,
@@
$\mathcal{L}_{RE}^{(i)}$ $=$ $-E_{\mathfrak{q}_{\theta_{e}}(\mathbf{z}|\mathbf{X}^{(i)})}\left[\log\mathfrak{p}_{\theta_{d}}(\mathbf{x}^{(i)}|\mathbf{z}^{\prime})\right],\;\;\mathcal{L}_{KL}^{(i)}=\beta\mathcal{D}_{KL}\left(\mathfrak{q}_{\theta_{e}}(\mathbf{z}|\mathbf{X}^{(i)})\,\|\,\tilde{\mathfrak{p}}_{\theta_{p}}(\mathbf{z})\right),$
@@
The distribution for the probabilistic encoding map is denoted by $\mathfrak{q}_{\theta_{e}}$, the distribution of the probabilistic decoding map by $\mathfrak{p}_{\theta_{d}}$, and the prior distribution by $\tilde{\mathfrak{p}}_{\theta_{d}}$. The prior $p(\mathbf{z})=\tilde{\mathfrak{p}}_{\theta_{p}}$ is used for regularizing the organization of the latent-space representation $\mathbf{z}$. Throughout, we use a Gaussian prior of the form $\tilde{\mathfrak{p}}_{\theta_{p}}(\mathbf{z})=\mathcal{N}(\mathbf{z};\boldsymbol{\mu}_{p},\boldsymbol{\Sigma}_{p})$ with mean zero $\boldsymbol{\mu}_{p}=0$ and diagonal covariance $\boldsymbol{\Sigma}_{p}=\sigma_{p}^{2}\mathcal{I}$. We remark other priors also can be considered that are further tailored to the task. The model is trained using batches of $m$ observation samples $\{(\mathbf{X}^{(i)},\mathbf{x}^{(i)})\}_{i=1}^{m}$. For a dynamical system with a collection of sampled trajectories $\{u^{k}(s)\}_{k}$, we have that $\mathbf{X}^{(i)}$ samples the initial state $u^{k_{i}}(t_{i})$ and $\mathbf{x}^{(i)}$ samples the future state $u^{k_{i}}(t_{i}+\tau)$. The $\tau$ gives the time-scale over which the prediction is made. The $k_{i}$ gives the trajectory index $k$ and the $t_{i}$ the sampled initial time $t$ from the trajectory.
The loss function in equation 8 provides a regularized form of MLE. The three terms comprising $\mathcal{L}^{B}$ can be interpreted as follows (i) $\mathcal{L}_{RR}$ for the log likelihood of reconstructing samples, (ii) $\mathcal{L}_{RE}$ is the log likelihood of correctly predicting samples after a single time step, and (iii) $\mathcal{L}_{KL}$ is a regularization term based on the Kullback-Leibler (KL) Divergence between the encoding distribution and latent space prior distribution.
@@
$\begin{array}[]{llllll}\mathbf{z}&\sim&\mathfrak{q}_{\theta_{e}}(\mathbf{z}|\mathbf{X}),&\mathbf{z}&=&\boldsymbol{\mu}_{e}(\mathbf{X};\theta_{e})+\boldsymbol{\eta}_{e}\\ \mathbf{z}^{\prime}&\sim&\mathfrak{p}_{\theta_{\ell}}(\mathbf{z}^{\prime}|\mathbf{z}),&\mathbf{z}^{\prime}&=&f_{\theta_{\ell}}(\mathbf{z})\\ \mathbf{x}&\sim&\mathfrak{p}_{\theta_{d}}(\mathbf{x}|\mathbf{z}^{\prime}),&\mathbf{x}&=&\boldsymbol{\mu}^{\prime}_{d}(\mathbf{z}^{\prime};\theta_{d})+\boldsymbol{\eta}_{d}^{\prime}\\ \mathbf{X}&\sim&\mathfrak{p}_{\theta_{d}}(\mathbf{X}|\mathbf{z}),&\mathbf{X}&=&\boldsymbol{\mu}_{d}(\mathbf{z};\theta_{d})+\boldsymbol{\eta}_{d}.\end{array}$ (16)